OpenStudy (anonymous):

Quick help: Explain how to solve the following system of equations. What is the solution to the system? 2x+2y+z=-5 3x+4y+2z=0 x+3y+2z=1

OpenStudy (solomonzelman):

I would use matrix

OpenStudy (anonymous):

I used matrix a wile ago I cannot remember how use it :/

OpenStudy (appleduardo):

check thi out and follow that pattern https://www.youtube.com/watch?v=TtxVGMWXMSE

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}? & ? & ? & ?\\ ? & ? & ? & ? \\ ? & ? & ? &?\end{matrix}\right] }\) lets start filling in numbers

OpenStudy (appleduardo):

this*

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 3 & 4 & 2 &~~~~~~ 0 \\ 1 & 3 & 2 &~~~~~~1\end{matrix}\right] }\)

OpenStudy (solomonzelman):

like this

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

you need to it to: \(\large\color{black}{ \left[\begin{matrix}1 & 0 & 0 &~~~~~~ x\\ 0 & 1 & 0 &~~~~~~ y \\ 0 & 0 & 1 &~~~~~~z\end{matrix}\right] }\)

OpenStudy (anonymous):

woah how'd you get that

OpenStudy (solomonzelman):

so lets try to zero out the 1st row 2nd column

OpenStudy (solomonzelman):

I would multiply the 3rd row, times -3, and add it to the second row.

OpenStudy (anonymous):

wait. How did you get the above matrix

OpenStudy (solomonzelman):

well, x y and z in the last column are just going to be the numbers that they (x y z) are equal to

OpenStudy (anonymous):

yeah, that part makes sense

OpenStudy (anonymous):

How did you get with the 1s and 0s

OpenStudy (solomonzelman):

that is what the set up should be at the end

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

so how would I get there?

OpenStudy (solomonzelman):

step by step, lol

OpenStudy (anonymous):

ok...

OpenStudy (solomonzelman):

Now, \(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 3 & 4 & 2 &~~~~~~ 0 \\ 1 & 3 & 2 &~~~~~~1\end{matrix}\right] }\) \(\large\color{black}{ \left[\begin{matrix} 1 & 3 & 2 &~~~~~~1\end{matrix}\right] }\) times -3, \(\large\color{black}{ \left[\begin{matrix} -3 & -9 & -6 &~~~~~~-3\end{matrix}\right] }\) adding to the 2nd row, \(\large\color{black}{ ~~~\left[\begin{matrix} -3 & -9 & -6 &~~~~~~-3\end{matrix}\right] }\) \(\large\color{black}{^+ \left[\begin{matrix}3 &~~~~~ 4 & ~~~~2 &~~~~~~~~~~~ 0\end{matrix}\right] }\)

OpenStudy (solomonzelman):

what will your second row become?

OpenStudy (anonymous):

Is every row times -3? always?

OpenStudy (solomonzelman):

I multiplied only the last row times -3, so that I can add (I am allowed to do this)

OpenStudy (solomonzelman):

what will the indicated sum (for second row) be?

OpenStudy (anonymous):

So am I multiplying row 2 and 1 by -3 as well?

OpenStudy (anonymous):

or is that -2 then -1?

OpenStudy (solomonzelman):

right now, you have just multiplied the last (3rd) row by -3, so that when you add it to the second row, the first number in second row becomes a zero.

OpenStudy (anonymous):

ok... so [0 -5 -4 -3]

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & -5 & -4 &~~~~~~ -3 \\ 1 & 3 & 2 &~~~~~~1\end{matrix}\right] }\)

OpenStudy (solomonzelman):

now, we will multiply the last row times -2, and leave it (for now) like this.

OpenStudy (anonymous):

then i do the sam eto the first [2 -3 -3 -8]

OpenStudy (solomonzelman):

what did you do?

OpenStudy (anonymous):

I did the second row to the first is that wrong?

OpenStudy (solomonzelman):

wrong direction, it is

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & -5 & -4 &~~~~~~ -3 \\ 1 & 3 & 2 &~~~~~~1\end{matrix}\right] }\) multiply the second row times -2, what do you get?

OpenStudy (solomonzelman):

I mean 3rd

OpenStudy (solomonzelman):

3rd row times -2

OpenStudy (anonymous):

[-2 -6 -4 -2]

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & -5 & -4 &~~~~~~ -3 \\ -2 & -6 & -4 &~~~~~~-2\end{matrix}\right] }\) yes.

OpenStudy (solomonzelman):

now add, the 1st row to the 3rd row. What will your third row be?

OpenStudy (anonymous):

[0 -4 -3 -7]

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & -5 & -4 &~~~~~~ -3 \\ 0 & -4 & -3 &~~~~~~-7\end{matrix}\right] }\)

OpenStudy (solomonzelman):

multiply the 2nd row times -1.

OpenStudy (anonymous):

Yeah i think i've got it. Could you help me with more questions?

OpenStudy (anonymous):

I'll open a new thingy though

OpenStudy (solomonzelman):

lets finish this, please. I don't want you to leave with an incorrect answer.

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & -5 & -4 &~~~~~~ -3 \\ 0 & -4 & -3 &~~~~~~-7\end{matrix}\right] }\) 2nd row times -1. \(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & 5 & 4 &~~~~~~ 3 \\ 0 & -4 & -3 &~~~~~~-7\end{matrix}\right] }\)

OpenStudy (solomonzelman):

add the 3rd row to the 2nd.

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 2 & 1 &~~~~~~ -5\\ 0 & 1 & 1 &~~~~~~ -4 \\ 0 & -4 & -3 &~~~~~~-7\end{matrix}\right] }\)

OpenStudy (anonymous):

i cant post something for some reason

OpenStudy (solomonzelman):

this will be quick, trust me a couple of minutes.

OpenStudy (anonymous):

ok, it isn't letting me post numbers

OpenStudy (solomonzelman):

refresh...

OpenStudy (anonymous):

I did lets see if this will work.

OpenStudy (anonymous):

aha ok so what do I need to be dong right now

OpenStudy (solomonzelman):

multiply the second row times -2, and add it to the 1st row. [ 0 - 2 -2 8] + [ 2 2 1 -5]

OpenStudy (solomonzelman):

what will your first row be?

OpenStudy (anonymous):

uh what is the stuff abover saying to multiply the second row

OpenStudy (solomonzelman):

I multiplied the second row times -2, and got [0 -2 -2 8] and then I am adding it to the 1st row, to make the first row better.

OpenStudy (solomonzelman):

[ 0 - 2 -2 8] + [ 2 2 1 -5] ------------------- ???

OpenStudy (anonymous):

oh ok so it will be 2 0 -1 3

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so we get: \(\large\color{black}{ \left[\begin{matrix}2 & 0 & -1 &~~~~~~ 3\\ 0 & 1 & 1 &~~~~~~ -4 \\ 0 & -4 & -3 &~~~~~~-7\end{matrix}\right] }\)

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

see how I am broaching the needed setup?

OpenStudy (anonymous):

yea?

OpenStudy (solomonzelman):

multiply the 2nd row times 4. add it to the last row. what will your last row be?

OpenStudy (anonymous):

4 or -4

OpenStudy (solomonzelman):

this time, only times positive 4

OpenStudy (anonymous):

0 4 4 -16

OpenStudy (solomonzelman):

yes, add this to the last row. what will your last row be?

OpenStudy (anonymous):

0 0 1 -23

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 0 & -1 &~~~~~~ 3\\ 0 & 1 & 1 &~~~~~~ -4 \\ 0 & 0 & 1 &~~~~~~-23\end{matrix}\right] }\)

OpenStudy (solomonzelman):

well, yes we have that z=-23. but we need the x and y too

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

add 3rd row to the 1st row.

OpenStudy (solomonzelman):

your 1st row becomes?

OpenStudy (anonymous):

2 0 0 -20

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

k

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}2 & 0 & 0 &~~~~~~ -20\\ 0 & 1 & 1 &~~~~~~ -4 \\ 0 & 0 & 1 &~~~~~~-23\end{matrix}\right] }\)

OpenStudy (solomonzelman):

divide first row by 2, you get?

OpenStudy (anonymous):

1 0 0 -10

OpenStudy (solomonzelman):

\(\large\color{black}{ \left[\begin{matrix}1 & 0 & 0 &~~~~~~ -10\\ 0 & 1 & 1 &~~~~~~ -4 \\ 0 & 0 & 1 &~~~~~~-23\end{matrix}\right] }\)

OpenStudy (solomonzelman):

multiply last row times -1, and add it to the 2nd row. you second row becomes?

OpenStudy (anonymous):

0 0 -1 23

OpenStudy (solomonzelman):

yes you multiplied times -1 correctly.

OpenStudy (solomonzelman):

now add this to the 2nd row

OpenStudy (anonymous):

0 1 0 19

OpenStudy (solomonzelman):

yes, so we get: \(\large\color{black}{ \left[\begin{matrix}1 & 0 & 0 &~~~~~~ -10\\ 0 & 1 & 0 &~~~~~~ 19\\ 0 & 0 & 1 &~~~~~~-23\end{matrix}\right] }\)

OpenStudy (solomonzelman):

clarifications?

OpenStudy (anonymous):

Cool thank you so much!! Nope! I'll start a new question in just a minute!

OpenStudy (solomonzelman):

I have to go light my Hannukah candels, see you:)